OpenStudy (calculusxy):

Chad poured a sidewalk that is 42 feet long. He is going to start at the beginning and make a cut every 20 inches. How many cuts will Chad make?

OpenStudy (johnweldon1993):

\[\large 42Feet \times \frac{12 Inches}{1 Foot} = 504 inches\] \[\large \frac{504}{20} = \space ?\]

OpenStudy (calculusxy):

So I did divide 504 by 20 but got 25 remainder 4. I don't know what to do with the remainder though. Should I remain it as 25 cuts or 26 cuts?

OpenStudy (johnweldon1993):

It would remain 25 cuts...just drop the remainder otherwise he wouldn't be sticking true to his "every 20 inches" rule

OpenStudy (calculusxy):

Okay thanks

OpenStudy (johnweldon1993):

No problem!